# Entanglement-breaking channels are a quantum memory resource

### v1

@@ Appendix D Numerical examples violating the classical the bound
$\begin{split}&\left|e_{1}\right\rangle=\begin{bmatrix}1\\ 0\\ 0\\ 0\end{bmatrix},\quad\left|e_{2}\right\rangle=\begin{bmatrix}0\\ \ -0.18168-0.03287i\\ -0.33582-0.01918i\\ 0.83859\end{bmatrix},\quad\left|e_{3}\right\rangle=\ \begin{bmatrix}0\\ 0.06342\,+0.62645i\\ -0.36075+0.0699i\\ -0.32475\end{bmatrix},\quad\ \left|e_{4}\right\rangle=\begin{bmatrix}0\\ -0.40607-0.42989i\\ -0.28345-0.24945i\\ -0.4251\end{bmatrix}\ ,\quad\left|e_{5}\right\rangle=\begin{bmatrix}0\\ -0.3906+0.2592i\\ 0.78055\\ 0.05633\,+0.08616i\ \end{bmatrix},\\ &\left|\phi_{1}\right\rangle=\begin{bmatrix}0\\ -0.25627-0.10917i\\ -0.29951-0.06935i\\ 0.90988\ \end{bmatrix},\quad\left|\phi_{2}\right\rangle=\begin{bmatrix}0\\ 0.78968\\ 0.17386\,+0.37575i\\ \ -0.02481+0.45209i\end{bmatrix},\quad\left|\phi_{3}\right\rangle=\begin{bmatrix}0\\ 0.73235\\ 0.46844\,\ -0.13164i\\ 0.35015\,-0.32293i\end{bmatrix},\quad\left|\phi_{4}\right\rangle=\begin{bmatrix}0\\ \ -0.34778+0.40375i\\ 0.82555\\ 0.14552\,+0.11544i\end{bmatrix},\quad\left|\phi_{5}\right\rangle=\ \begin{bmatrix}1\\ 0\\ 0\\ 0\end{bmatrix},\\ &K_{0}=\mathrm{diag}(0,1,1,1),\qquad K_{1}=\mathrm{diag}(1,0,0,0).\end{split}$ (15)
For $L=3$ and $d=2$, we were unable to find any violation of the classical bound, strongly suggesting EB-based temporal non-classicality only emerges beyond two-level systems. Furthermore, since $\operatorname{DC}(\bm{a})\leq L$ for any sequence of length $L$, either a qubit or a bit are sufficient to produce any correlations on two time steps. It is therefore essential to also look beyond two-time steps in order to reveal fundamental differences between classical and quantum systems in this sequential measurement scenario.

### v2

@@ Appendix D Numerical examples violating the classical bound
$\begin{split}&\left|e_{1}\right\rangle=\begin{bmatrix}1\\ 0\\ 0\\ 0\end{bmatrix},\quad\left|e_{2}\right\rangle=\begin{bmatrix}0\\ \ -0.18168-0.03287i\\ -0.33582-0.01918i\\ 0.83859\end{bmatrix},\quad\left|e_{3}\right\rangle=\ \begin{bmatrix}0\\ 0.06342\,+0.62645i\\ -0.36075+0.0699i\\ -0.32475\end{bmatrix},\quad\ \left|e_{4}\right\rangle=\begin{bmatrix}0\\ -0.40607-0.42989i\\ -0.28345-0.24945i\\ -0.4251\end{bmatrix},\\ &\left|e_{5}\right\rangle=\begin{bmatrix}0\\ -0.3906+0.2592i\\ 0.78055\\ 0.05633\,+0.08616i\ \end{bmatrix},\quad\left|\phi_{1}\right\rangle=\begin{bmatrix}0\\ -0.25627-0.10917i\\ -0.29951-0.06935i\\ 0.90988\ \end{bmatrix},\quad\left|\phi_{2}\right\rangle=\begin{bmatrix}0\\ 0.78968\\ 0.17386\,+0.37575i\\ \ -0.02481+0.45209i\end{bmatrix},\\ &\quad\left|\phi_{3}\right\rangle=\begin{bmatrix}0\\ 0.73235\\ 0.46844\,\ -0.13164i\\ 0.35015\,-0.32293i\end{bmatrix},\quad\left|\phi_{4}\right\rangle=\begin{bmatrix}0\\ \ -0.34778+0.40375i\\ 0.82555\\ 0.14552\,+0.11544i\end{bmatrix},\quad\left|\phi_{5}\right\rangle=\ \begin{bmatrix}1\\ 0\\ 0\\ 0\end{bmatrix},\\ &K_{0}=\mathrm{diag}(0,1,1,1),\qquad K_{1}=\mathrm{diag}(1,0,0,0).\end{split}$ (16)